### v2

# Tilus: A Virtual Machine for Arbitrary Low-Precision GPGPU Computation in LLM Serving

Abstract.
Serving Large Language Models (LLMs) is critical for AI-powered applications but demands substantial computational resources, particularly in memory bandwidth and computational throughput. Low-precision computation has emerged as a key technique to improve efficiency while reducing resource consumption. Existing approaches for generating low-precision kernels are limited to weight bit widths that are powers of two and suffer from suboptimal performance due to high-level GPU programming abstractions. These abstractions restrict critical optimizations, such as fine-grained register management and optimized memory access patterns, which are essential for efficient low-precision computations. In this paper, we introduce a virtual machine (VM) designed for General-Purpose GPU (GPGPU) computing, enabling support for low-precision data types with arbitrary bit widths while maintaining GPU programmability. The proposed VM features a thread-block-level programming model, a hierarchical memory space, a novel algebraic layout system, and extensive support for diverse low-precision data types. VM programs are compiled into highly efficient GPU programs with automatic vectorization and instruction selection. Extensive experiments demonstrate that our VM efficiently supports a full spectrum of low-precision data types, and outperforms state-of-the-art low-precision kernels on their supported types. Compared to existing compilers like Triton and Ladder, as well as hand-optimized kernels such as QuantLLM and Marlin, our VM achieves performance improvements of 1.75×1.75$\times$1.75 ×, 2.61×2.61$\times$2.61 ×, 1.29×1.29$\times$1.29 × and 1.03×1.03$\times$1.03 ×, respectively.
1. Introduction
The development of Large Language Models (LLMs) has revolutionized natural language processing tasks, enabling advanced capabilities in areas such as text generation (Brown et al., 2020), summarization (Liu and Lapata, 2019), translation (Workshop et al., 2022), and conversational AI (OpenAI, 2024). However, serving LLMs poses substantial computational challenges due to the large model sizes and high computational demands. Efficient LLM serving demands innovative computational strategies to manage latency and power consumption constraints. As such, optimizing LLM inference has become a priority in both industry and research to reduce latency and increase throughput of LLM serving.
Quantization (Frantar et al., 2022; Lin et al., 2024; Ashkboos et al., 2024; Chee et al., 2023; Tseng et al., 2024; Liu et al., 2025) has emerged as a leading method for enhancing the efficiency of LLM serving. By reducing the bit-width of model parameters and activations, quantization reduces weight storage, DRAM bandwidth usage, and achieves faster computation. For instance, A16W4 quantization (16-bit activation and 4-bit weight) reduces the DRAM consumption and throughput by 4×4$\times$4 × compared with A16W16 scheme, thus reducing the time to generate a token by about 4×4$\times$4 × (Frantar et al., 2024). However, the state-of-the-art 4-bit quantization methods (Ashkboos et al., 2024; Chee et al., 2023; Liu et al., 2025) still suffer from non-negligible accuracy degradation. While using 5- to 7-bit quantization could mitigate this accuracy loss (Agrawal et al., 2024; Xia et al., 2024), the lack of efficient GPU kernels for these bit widths has hindered their adoption. Generating optimized kernels for hardware-unfriendly bit widths (e.g., 3, 5, 6, and 7) remains an open problem.
Existing methods for generating computation kernels fall into two main categories: manually written kernels (Frantar et al., 2024; Xia et al., 2024) and compiler-generated kernels (Chen et al., 2018; Wang et al., 2024; Tillet et al., 2019). While manually written kernels are highly optimized for specific hardware, they are time-consuming and error-prone to develop, and difficult to generalize to new architectures and evolving quantization methods. For example, QuantLLM (Xia et al., 2024) only supports floating-point quantization for 5- and 6-bit data types but lacks support for sub-channel quantization granularity. Similarly, Marlin (Frantar et al., 2024) is limited to 4-bit signed integer quantization and does not support Hopper GPUs (Luo et al., 2025).
To address these limitations, compiler-based approaches (Chen et al., 2018; Wang et al., 2024; Tillet et al., 2019) have been proposed to automate kernel generation. Among them, Triton (Tillet et al., 2019) simplifies the GPGPU programming through a tile-based model. A Triton program defines the computation of a thread block over tensor tiles. However, Triton lacks built-in support for low-precision data types, requiring users to implement low-level bitwise operations manually. Additionally, it does not expose the full GPU memory hierarchy, limiting optimization opportunities for low-precision LLM inference. Ladder (Wang et al., 2024), on the other hand, extends TVM’s scheduling system (Chen et al., 2018) to support low-precision computation but is restricted to data types with bit widths that are powers of two. Moreover, its primitives cannot express crucial optimizations such as software pipelining (Huang et al., 2023), leading to suboptimal performance, particularly for batch sizes greater than one during LLM decoding with continuous batching (Yu et al., 2022) enabled.
To address these challenges, we propose Tilus, a GPGPU virtual machine with dedicated support for low-precision computation. Tilus abstracts GPU program execution into thread-block-level instructions, simplifying GPGPU programming while exposing hierarchical memory spaces for fine-grained manipulation of sub-tensors in on-chip memory. This dual approach enables efficient handling of arbitrary-precision data types while reducing the complexity of GPU programming. To achieve these goals, Tilus introduces: (1) an algebraic layout system that specifies how tensor elements within a tile are distributed across GPU threads. This layout system enables flexible reinterpretation of low-precision tiles in registers as tiles with hardware-friendly data types for efficient processing; (2) a thread-block-level programming model with fine-grained memory management, providing explicit control over data movement, placement, and computation across different levels of the GPU memory hierarchy; and comprehensive support for (3) arbitrary low-precision data types including signed integers, unsigned integers, and floating-point numbers with 1 to 8 bits.
Extensive experiments show that Tilus extends the spectrum of efficient low-precision kernels to support arbitrary bit widths (1-8) and data type kinds (e.g., integer and floating numbers). At the same time, Tilus outperforms the state-of-the-art compiler Triton (Tillet et al., 2019), Ladder (Wang et al., 2024), and hand-crafted kernels from QuantLLM (Xia et al., 2024) and Marlin (Frantar et al., 2024) by 1.75×1.75$\times$1.75 ×, 2.61×2.61$\times$2.61 ×, 1.29×1.29$\times$1.29 × and 1.03×1.03$\times$1.03 × on their supported low-precision kernels, respectively.
We summarize our key contributions as follows:
  • We propose a GPGPU virtual machine with dedicated support for low-precision computation, addressing the critical coverage (e.g., support 5 to 7 bits quantization) and performance gap in existing approaches.
  • Inside the virtual machine, we introduce a novel layout system, a thread-block-level programming model with hierarchical memory space, and support low-precision data types with arbitrary bit widths.
  • Through extensive evaluation, we demonstrate that Tilus generates a full spectrum of highly efficient low-precision kernels, outperforming state-of-the-art solutions across their supported kernels by up to 2.6×2.6$\times$2.6 ×.
2. Background
2.1. LLM Serving and Quantization
In large language model (LLM) serving, inference consists of two stages: prefill and decode. The prefill stage processes the input prompt to establish context, while the decode stage iteratively generates output tokens based on prior tokens. Among all layers of LLMs, matrix multiplications dominate computation time and memory consumption, making their optimization crucial for efficient LLM serving. Quantization (Dettmers et al., 2024; Frantar et al., 2022) improves their efficiency by reducing model weights and activations to lower-precision formats, such as 8-bit or 4-bit integers. It reduces memory usage, bandwidth requirements, and inference latency while trying to maintain model accuracy. While 4-bit quantization provides significant computational savings, state-of-the-art methods (Ashkboos et al., 2024; Chee et al., 2023; Liu et al., 2025) still suffer from accuracy degradation. Increasing precision to 5-bit, 6-bit, or 7-bit quantization (Agrawal et al., 2024; Xia et al., 2024) can help preserve accuracy while maintaining efficiency, but these bit widths lack optimized GPU support, limiting their adoption. Current GPU architectures and software stacks primarily optimize for power-of-two bit widths (e.g., 4-bit and 8-bit), making arbitrary bit widths computationally inefficient. However, demand for flexible quantization is growing, as 4-bit can be too aggressive for some models while 8-bit wastes resources. Supporting a broader spectrum of bit widths enables better accuracy-efficiency trade-offs in LLM serving, driving the need for new kernel generation techniques that can efficiently handle non-standard low-precision formats (e.g., those with 3, 5, 6, 7 bit widths) on modern GPUs.
2.2. The GPGPU Programming
General-Purpose GPU (GPGPU) programming enables parallel computation by organizing tasks within a structured execution and memory hierarchy (Nickolls and Dally, 2010). The execution hierarchy begins with the thread, the smallest unit of execution, which performs instructions independently with its own registers and local memory. Threads are grouped into thread blocks, which enable data sharing through shared memory and support synchronized execution. A grid consists of multiple independent thread blocks, enabling large-scale parallelism by organizing thousands or millions of threads. The GPU memory hierarchy comprises registers, shared memory, and global memory. Registers provide the fastest and thread-private storage. Shared memory is accessible by all threads within a thread block and faster than global memory. Global memory is accessible across the entire grid with high latency. This structure allows for highly efficient parallel execution by leveraging both the execution and memory hierarchies.
2.3. The GPGPU Languages and Compilers
2.3.1. GPGPU Virtual Machine and Languages
GPGPU programming involves various languages and compilers that balance hardware abstraction with control. Low-level languages like SASS (NVIDIA Corporation, 2024c) and CDNA3 (AMD Corporation, 2024a) offer direct hardware access for fine-grained optimizations but require deep architectural knowledge. Slightly higher in abstraction, NVIDIA’s PTX (NVIDIA Corporation, 2024b) serves as an intermediate representation, linking high-level languages such as CUDA (NVIDIA Corporation, 2024a) to GPU-specific instructions while preserving optimization flexibility. High-level languages like CUDA (NVIDIA Corporation, 2024a) and HIP (AMD Corporation, 2024b) simplify programming by extending C programming language. Despite these languages, GPGPU programming remains complex. It is constrained by hardware-specific memory and computation hierarchies and requires workload-specific optimizations. To address these challenges, researchers have introduced higher-level languages and compilers, classified into two categories: procedure-oriented compilers, which simplify programming through abstractions beyond CUDA (NVIDIA Corporation, 2024a), and schedule-oriented compilers, which optimize computation-hardware mappings via declarative scheduling primitives.
2.3.2. Procedure-Oriented Compilers
This type of compilers (Lattner et al., 2021; Hagedorn et al., 2023; Ding et al., 2023; Tillet et al., 2019) enables programmers to write kernels directly, offering abstractions to simplify the process. Triton (Tillet et al., 2019), for instance, introduces the tile programming model, where thread block behavior is defined programmatically, and tiles replace scalars as the basic data type. This approach combines programming simplicity with high-performance kernel generation, making Triton widely adopted. However, Triton lacks native support for low precision data types like uint4. Handling these types requires manually unpacking sub-byte data from larger storage types (e.g., uint32(Hoque et al., 2024). Additionally, Triton does not expose the GPU memory hierarchy, limiting programmers’ control over data loading and memory scope usage, which complicates performance optimization for low-precision kernels. These limitations result in inefficient low-precision kernel execution. Figure 1(a) illustrates the inefficiencies in Triton-generated low-precision kernels, using a uint4 weight loading pipeline as an example. The process includes four steps: 1 weights are asynchronously copied from global memory to shared memory using pipelined cp.async instructions (Huang et al., 2023); 2 shared memory data is loaded into registers; 3 unpacking and casting operations are performed; and 4 the register tensor layout is converted to meet the requirements of tensor core instructions. Among these, step 4 is a major bottleneck due to the reliance on shared memory for layout conversion, which incurs significant overhead.
2.3.3. Schedule-Oriented Compilers
Schedule-oriented compilers decouple computation from scheduling to optimize the computation-to-hardware mappings. Halide (Ragan-Kelley et al., 2013) pioneered this approach, which was later extended by TVM (Chen et al., 2018) and subsequent works (Vaswani et al., 2017; Zheng et al., 2020a, 2022a; Shao et al., 2024; Huang et al., 2023; Feng et al., 2023; Wang et al., 2024) in the domain of deep learning. Among them, Ladder (Wang et al., 2024) is the first one to support low-precision computation by introducing dedicated primitives to pack low-precision data (e.g., 4-bit integers) into larger types (e.g., 8-bit integers). However, Ladder (Wang et al., 2024) has two limitations. First, it cannot handle non-power-of-two bit widths efficiently due to type-level packing, packing low-precision types into storage types. Second, its primitive-style scheduling prevents optimizations like software pipelining (Huang et al., 2023), resulting in suboptimal performance. Figure 1 (b) illustrates the weight loading process in Ladder’s low-precision kernels. This process includes 1 loading weights from global memory to registers without pipelining; 2 vectorized casting; 3 storing the cast results in shared memory; and finally 4 using the ldmatrix instruction to load weights from shared memory to registers for subsequent tensor core operations. The lack of pipelining between weight loading and computation significantly hinders performance.
[…]
Weight loading pipeline illustration
Figure 1. The weight loading pipeline of Triton, Ladder, and our approach. The tensors could be in either global memory (GMEM), shared memory (SMEM), and registers (REGS).
Figure 2. This figure provides a concrete example of how the GPGPU virtual machine is used to implement low-precision matrix multiplication (FP16 $\times$ INT6). Figure (a) illustrates the virtual machine program, highlighting key features such as the algebraic layout system (Section 4), thread-block-level instructions (Section 5), and efficient low-precision data handling. Figure (b) illustrates the kernel’s data flow, emphasizing tensor movement across the memory hierarchy and intermediate operations such as tensor reinterpretation and type casting. A similar weight-loading strategy can be applied to arbitrary type widths (Section 6). Finally, Figure (c) demonstrates register tensor reinterpretation, showing how tensors with compatible bit distributions across threads (e.g., 24 bits per thread) can be efficiently reinterpreted into different data types and layouts.
[…]
matmul example.
3. System Overview
3.1. Key Ideas
Our work introduces a novel GPGPU virtual machine (VM) specifically designed to overcome the challenges of programming efficient low-precision deep learning kernels. The VM natively supports low-precision data types with arbitrary bit widths ranging from 1 to 8, enabling efficient weight loading and computation. Figure 1 (c) shows the VM’s weight loading pipeline, using uint4 as an example. It begins with 1 a pipelined asynchronous memory copy from global memory to shared memory, followed by 2 loading the register tensor from shared memory. Next, it 3 reinterprets the register tensor into a different data type and layout at no cost, and then 4 performs vectorized casting. This pipeline achieves superior efficiency compared to the other methods in Figure 1, as it eliminates layout conversion (unlike Triton (Tillet et al., 2019)) and incorporates pipelining (unlike Ladder (Wang et al., 2024)). More importantly, our pipeline is versatile, making our work the first to seamlessly support arbitrary low-precision data types with bit widths ranging from 1 bit to 8 bits.
To achieve this efficiency, our design is built on several key ideas. A GPGPU Virtual Machine for Flexible Optimization: our decision to implement a GPGPU VM stems from the need for greater flexibility in GPU programming. Unlike standard loop-based transformations, our VM allows programmers to directly implement and fine-tune optimizations that go beyond traditional loop transformations. This flexibility is critical for low-precision computations, where fine-grained control over execution strategies can lead to significant performance gains. A Thread-Block-Level Programming Model with Hierarchical Memory Spaces: our VM explicitly exposes the GPU memory hierarchy, including registers, shared memory, and global memory, that are abstracted away in existing solutions like Triton (Tillet et al., 2019). By granting programmers fine-grained control over data placement and movement, our approach enables memory pipelining and eliminates unnecessary layout conversions, as shown in Figure 1. An Algebraic Layout System: we introduce an algebraic layout system that precisely defines how elements within a register tensor are distributed among threads. This structured representation simplifies the construction, analysis, and interpretation of tensor layouts. Notably, it enables seamless reinterpretation of low-precision register tensors into standard data types, as demonstrated in Step 3 of Figure 1(c). Native Support for Arbitrary Low-Precision Data Types: our VM provides built-in support for a wide range of low-precision data types, including both signed and unsigned integers and floating-point numbers with bit widths from 1 to 8. Supported types include int2 to int8, uint1 to uint8, and float3 to float8, with arbitrary exponent and mantissa distribution for floating-point types. These innovations collectively enhance the programmability, efficiency, and flexibility of low-precision kernel development on modern GPUs. We chose not to extend Triton (Tillet et al., 2019) because its programming model inherently abstracts away tensor layouts, making it incompatible with our approach of explicit layout control. Similarly, Ladder (Wang et al., 2024) relies on type-level packing, whereas Tilus employs tile-level reinterpretation, making the two fundamentally incompatible. The next section presents an example of low-precision matrix multiplication within our virtual machine.
3.2. An Example of Virtual Machine Program
Figure 2 illustrates a low-precision matrix multiplication in the virtual machine. Matrix multiplication is defined as $C_{M,N}=A_{M,K}\times B_{K,N}$, where $A$ and $B$ are of type float16 (a 16-bit floating-point number (IEEE Standards Association, 2019)) and int6 (a 6-bit signed integer), respectively. The kernel performs matrix multiplication with given M, N, and K, where each thread block computes a BM $\times$ BN tile of the C matrix (Line 1). Therefore, a grid of (M / BM, N / BN) thread blocks must be launched (Line 2). Inside the kernel, the BlockIndices instruction retrieves the thread block indices bi and bj (Line 3), which determine the offset (bi * BM, bj * BN) for computing the corresponding C tile. Three tensor views are created for the input and output tensors in global memory by specifying their addresses and shapes (Line 4-6). Then, a register tensor of type f16[16, 8] is created with the following layout:
local(2, 1).spatial(8, 4).local(1, 2).
It distributes $16\times 8=128$ elements across 32 threads, with each thread storing 4 elements (Line 7). This layout is composed of three primitive layouts (Section 4) and aligns with the C matrix layout used by the mma.m16n8k16 tensor core instruction in PTX (NVIDIA Corporation, 2024b). The reduction loop over the k dimension (Line 8-13) repeatedly loads tiles of A and B from global memory into registers and performs matrix-multiply accumulate (mma). For each iteration, we first load a f16[16, 16] tile from global memory to register with a LoadGlobal instruction (Line 9). The layout of the loaded register tile is specified and required by the tensor core instruction. The offset parameter specifies the position of the loaded tile within the global tensor. Loading tensor B, which has data type int6, involves a more complex process, detailed in Section 6. We summarize the high-level ideas here. As a pre-processing step before launching the kernel, the weight tensor’s layout in global memory is transformed from i6[K, N] to u8[K / BK, N / BN, BK * BN * 6 / 8], enabling efficient loading via the LoadGlobal instruction (the ‘Change Layout’ step in Figure 2 (b)). Next, in the kernel, the transformed tile is loaded into a register tensor (Line 10) and then reinterpreted to a tensor with a different data type and layout (Line 11). This reinterpretation is valid because both tensors are stored across the same number of threads (32), with each thread holding exactly 24 bits (i.e., 3 $\times$ u8 or 4 $\times$ i6), as shown in Figure 2 (c). Following this, the i6 tensor is cast to an f16 tensor (Line 12), which is then fed to the tensor core to perform matrix-multiply accumulate (mma) (Line 13). Finally, the accumulation tensor is cast from f32 to f16 and stored in global memory (Line 14-15). For simplicity, this program does not utilize shared memory and omits optimizations such as software pipelining (Huang et al., 2023); additionally, each k-iteration performs only a single tensor core instruction (NVIDIA Corporation, 2024a). An optimized implementation can be found in Appendix B.
The following sections introduce the three core components of the proposed VM. Section 4 introduces an algebraic layout formulation to systematically define how the elements of a tile are stored in the registers among the block threads. Section 5 introduces the thread-block-level programming model with a hierarchical memory space exposed explicitly. Section 6 introduces the native support for arbitrary low-precision data types to address the growing demand for low-precision computation in deep learning workloads.
4. Algebriac Layout System
Figure 3. Layout of operand A in a Tensor Core instruction. The operand, with $16\times 8$ elements, is distributed across $32$ threads, with each thread storing four elements. The logical index of each element is determined by a layout function given the thread index t and the local element index i.
[…]
mma layout
The virtual machine exposes a hierarchical memory space with global memory, shared memory, and registers to programmers. We need a way to model the mapping between the logical index of a tensor element and the location of the corresponding element in memory for all three memory scopes. Such a mapping is usually called the layout of the tensor. Figure 3 illustrates an example of the layout used by a tensor core instruction: mma.m16n8k8.f32.f16.f16.f32 D, A, B, C. It performs the following computation: $D_{16,8}=A_{16,8}B_{8,8}+C_{16,8}$ where $A,B,C,D$ are tensors stored in thread registers and distributed across the $32$ threads in a warp. Since the elements are spread across different threads, we refer to this layout as a distributed layout (Tillet et al., 2019). Such a layout can be defined as a function $f$ that maps a thread index $t$ and a local index $i$ within that thread to the logical index $f(t,i)$ of the corresponding tensor element. For example, the layout in Figure 3 can be represented as:
$f(t,i)=(t/4+i/2\times 8,t\ \%\ 4\times 2+i\ \%\ 2)$
Here, $t$ ranges from 00 to $31$, and $i$ ranges from 00 to $3$. The function $f(t,i)$ represents the logical index of element $i$ in thread $t$. Since all threads in a thread block have access to shared and global memory, this formalization can also be used to describe their layouts under a single-thread assumption. That is, by setting $t=0$, we define $f(0,i)$ as the logical index of the element at address $i$ in shared or global memory.
4.1. Parameterized Primitive Layouts
Figure 4. Two types of primitive layouts: local and spatial. A local layout stores all tile elements within a single thread, whereas a spatial layout distributes them across multiple threads, with each thread holding only a single element.
[…]
primitive layouts
With the formal definition of layout, we introduce parameterized primitive layouts that serve as the fundamental building blocks of our layout algebra. Given a tile [^2] We use the terms tile and tensor interchangeably. with shape $(n_{1},n_{2})$, there are two primary ways to store it: (1) storing all $n_{1}n_{2}$ elements in a single thread, or (2) distributing all elements across $n_{1}n_{2}$ threads, where each thread holds a single element. We refer to the first type as local layouts, denoted as local(n1, n2), and the second as spatial layouts, denoted as spatial(n1, n2). This concept naturally extends to tiles with arbitrary dimensions. Figure 4 illustrates these two primitive layouts. The local(2, 3) layout maps the $i$-th local element of thread $t$ to the logical index $(i/3,i\ \%\ 3)$, while the spatial(2, 3) layout maps it to $(t/3,t\ \%\ 3)$. We observe that all complex layouts used in modern deep learning workloads and GPUs can be constructed using these two primitive layouts. In the next section, we introduce methods for composing these layouts to define more complex layouts.
4.2. Layout Composition
Figure 5. Complex layouts can be constructed through composition. In the figure, layout (c) is composed of layouts (a) and (b), while layout (e) is composed of layouts (d) and (a).
[…]
layout composition
The layouts used in modern deep learning workloads, as well as those defined by hardware instructions, typically exhibit a hierarchical structure. Consider layout (c) in Figure 5 as an example. This layout has shape $(4,6)$, storing $24$ elements across $6$ threads, with each thread holding four elements. We denote the four elements stored in each thread as $a_{0},a_{1},a_{2},a_{3}$. Comparing its first two rows with the last two, we observe a similar structure, except that the last two rows store elements in $a_{2}$ and $a_{3}$ instead of $a_{0}$ and $a_{1}$. To model this structural invariance, layout (c) can be viewed as a composition of layouts (a) and (b), where each element in layout (a) represents a tile with layout (b). In fact, layouts (a) and (b) can be combined to represent layout (c) as follows:
$c(t,i)$ $=a(t/6,i/2)\odot(2,6)+b(t\ \%\ 6,i\ \%\ 2),$
where $60\leq t<$ 6, $40\leq i<$ 4, $\odot$ denotes the element-wise product, and $(2,6)$ represents the shape of layout (c). This composition principle can be generalized. Given two layouts $f$ and $g$ with the same number of dimensions, we define their composition $h=f\circ g$ as
$h(t,i)=f(t/T_{g},i/N_{g})\odot S_{g}+g(t\ \%\ T_{g},i\ \%\ N_{g}),$
where $T_{g},N_{g},S_{g}$ represent the number of threads, the number of local elements per thread, and the shape of layout $g$, respectively. We can prove that composition is associative, meaning that for any three layouts $f$, $g$, and $h$, the equality $f\circ(g\circ h)=(f\circ g)\circ h$ holds. However, composition is not commutative, meaning that, in general, $f\circ g\neq g\circ f$. Spatial and local layouts follow a row-major ordering for threads and local elements, respectively. Using composition, we can construct their column-major counterparts, column_spatial(...) and column_repeat(...), as demonstrated by layout (e) in Figure 5. Returning to the tensor core instruction layout in Figure 3, it can be expressed as a composed layout local(2, 1).spatial(8, 4).local(1, 2). Using layout composition, we can also define its inverse operation. If $h=f\circ g$, we define $f=h/g$ as the result of dividing layout $h$ by layout $g$. For example, dividing local(2, 4) by local(1, 2) results in local(2, 2). In Appendix A, we formally define layouts and prove that under the composition operator, they form a semigroup with an identity element, making them a monoid.
5. Thread-Block-Level Programming Model
Modern GPU programming models, such as PTX (NVIDIA Corporation, 2024b) and CUDA (NVIDIA Corporation, 2024a), define operations at the thread level, following the Single-Instruction-Multiple-Thread (SIMT) paradigm (Aamodt et al., 2018). To simplify GPU programming, we adopt the thread-block-level programming model, defining operations at the granularity of thread blocks rather than individual threads. Additionally, building on the layout system introduced previously, we propose explicitly exposing the hierarchical memory structure in modern GPUs, enabling fine-grained memory control while reducing programming complexity. We refer to this model as Single-Instruction-Multiple-Block (SIMB).
5.1. State Space and Type System
The virtual machine supports three types of variables. Scalar variables store individual values, such as integers (e.g., int32) or floating-point numbers (e.g., float16). Pointer variables store memory addresses rather than direct data values. Tensor variables represent multi-dimensional arrays, with types that specify their shape, element type, memory scope, and layout. Tensors reside in different memory scopes, including global memory, shared memory, and registers. The tensor layout determines how high-dimensional tensor elements are mapped to linear memory. All variables in the virtual machine operate at the thread-block level, meaning that all threads within a block collaboratively maintain them.
5.2. Program Structure and Control Flow
Figure 6. Virtual machine program structure. A virtual machine program contains parameters and a body. The body is a list of control-flow statements or block-level instructions.
[…]
instruction
Table 1. The thread-block-level instruction set of the virtual machine. Each instruction specifies an operation applied to the entire thread block. Parameters enclosed in [...] are optional. Instructions that return a new register tensor also have an in-place variant, which writes the result to an existing register tensor using the out parameter instead of creating a new tensor.
Figure 6 illustrates the structure of a VM program. Each program consists of a program name, a grid shape, a list of parameters, and a program body. The grid shape is specified as a list of expressions enclosed in <...>, where each expression is either a positive integer or an integer expression based on the program parameters. If the grid shape contains parameter-based expressions, its dimensions are determined at runtime based on the program’s launch arguments. The program body consists of a sequence of statements, including if-else statements, range-based for-loops, and while-loops. Unlike other low-level virtual machines (NVIDIA Corporation, 2024b) or instruction set architectures (ISAs) (NVIDIA Corporation, 2024c), our virtual machine does not abstract control-flow statements into jump instructions. Instead, it retains high-level control structures to improve readability and ease of programming for human developers. In addition to control-flow statements, individual instructions can also serve as statements. Most functionalities provided by the virtual machine are implemented as instructions in the thread-block-level instruction set.
5.3. Thread-Block-Level Instruction Set
Each instruction in the VM’s instruction set operates at the thread-block level rather than the thread level. Table 1 shows a list of the instructions in the instruction set, with the signature of each instruction and a brief description of the instruction semantics. These instructions allocate tensors with specific data types, shapes, and layouts in designated memory spaces (e.g., global memory, shared memory, registers), transfer tensors between memory spaces, and perform computations or transformations on register tensors. Modern processors execute instructions out of order (Tomasulo, 1967), meaning that a subsequent instruction may begin execution before the current one completes, provided there are no dependencies between them. Similarly, the execution of instructions in our virtual machine exhibits this behavior: certain subsequent instructions may begin execution before the current instruction completes. Generally, this behavior does not pose significant issues. However, an exception occurs when two instructions access the same region of shared or global memory, and the second instruction depends on the completion of the first. In such cases, a Synchronize instruction must be inserted to ensure all preceding instructions complete before subsequent ones execute.
6. Arbitrary Low-Precision Data Types
Modern processors use bytes (8 bits each) as the smallest processing unit. As a result, standard data types in modern programming languages typically have bit widths that are multiples of 8. However, the high computational and memory demands of LLMs make low-precision data types (less than 8 bits) essential for reducing resource consumption. This section describes how the virtual machine efficiently supports low-precision data types.
6.1. Storage of Low-Precision Data
Figure 7. Compact storage and access of low-precision data. Figure (a) illustrates the use of the uint8 type to store low-precision data, where some elements may span two consecutive bytes. Figure (b–c) illustrate the implementation of loading and storing low-precision elements.
Since modern processors, including CPUs and GPUs, use bytes as the smallest unit for memory access and computation, we store low-precision data (fewer than 8 bits per element) compactly within bytes, as shown in Figure 7. Compact storage eliminates bit gaps between consecutive low-precision values, which may result in a single value spanning two uint8 entries (e.g., b[1] in Figure 7). Bitwise operations are employed to extract, manipulate, and store low-precision values within packed byte arrays. To load a low-precision value, we first extract relevant bits using bitwise AND, adjust their position with bitwise SHIFT operations, and finally combine separated parts using bitwise OR if the value spans multiple bytes. Similarly, to store a low-precision value, we first clear the target bit positions using a bitwise mask, then insert the new value using bitwise OR while preserving the other bits. Low-precision data is cast to standard data types before arithmetic computations and cast back afterward. While these methods enable support for arbitrary bit-width data types, they are often inefficient and serve only as a fallback mechanism. More efficient handling of low-precision data is necessary for LLM serving.
6.2. Efficient Low-Precision Support in LLMs
Figure 8. Program to rearrange tensor B with data type int6, used in the ”Change Layout” step of Figure 2 (b).
Low-precision kernels in LLMs typically follow two steps before computation: (1) loading weights into on-chip memory (registers or shared memory) from global memory, and (2) casting low-precision weights to high-precision (e.g., float16) followed by de-quantization. Efficient memory loading and casting are thus critical for performance.
[…]
Efficient Casting. After loading, weights must be cast from low-precision to high-precision (e.g., float16) for computation, especially if hardware lacks native support for the given low-precision format. We leverage target-specific instructions for efficient vectorized casting. On CUDA, we use the PRMT (permute bytes in a 32-bit register), LOP3 (arbitrary logical operation on three inputs), and bitwise instructions to execute casting with minimal overhead, as all operations are performed within registers and do not require any communication between threads.
7. Implementation
Tilus comprises five main components: a domain specific language (DSL) in Python, an intermediate representation (IR), optimization passes, a code generator, and a runtime system. The DSL enables developers to write Tilus programs in Python, which are then translated into the VM’s IR for further processing. Optimization passes refine the IR by eliminating redundancies and simplifying arithmetic expressions. The code generator translates the optimized IR into Hidet IR (Ding et al., 2023), a CUDA C-like intermediate representation. Afterwards, we apply the transformations from Section 6 to implement low-precision types with standard precision types while preserving original semantics. The final CUDA C code is generated from Hidet IR and compiled into a hardware binary using the nvcc compiler (NVIDIA Corporation, 2024a). The runtime system manages dynamically loaded binaries and provides the execution environment. The entire system consists of around 20K lines of Python and C++ code. Further details on the compilation pipeline can be found in Appendix C.
Figure 9. Speedup of low-precision kernels in Triton, QuantLLM, Ladder, and Tilus (Ours) compared against the standard half-precision kernel from cuBLAS. Benchmarked data types include uint8 (u8), f6e3m2 (f6), int4 (i4), uint4 (u4), uint2 (u2), and uint1 (u1). Each workload (BS-N-K) corresponds to a matrix multiplication in Llama-3.3-70B, with batch sizes 1 and 16.
8. Evaluation
8.1. Experimental Setup
Workloads. We benchmark three representative LLMs with varying model sizes: Gemma-2-9B (Team et al., 2024), QWen2.5-32B (Yang et al., 2024), and Llama-3.3-70B-Instruct (Grattafiori et al., 2024). Both the prefill and decode stages are evaluated. For operator-level analysis, we focus on matrix multiplication kernels extracted from these models. Tilus supports all kernels supported by Triton in principle, but we focus on quantized matmul in this work.
Baselines. We compare our approach, Tilus, against the vendor library cuBLAS (NVIDIA Corporation, 2023), state-of-the-art DL compilers Triton (Tillet et al., 2019) and Ladder (Wang et al., 2024), and hand-crafted kernels QuantLLM (Xia et al., 2024) and Marlin (Frantar et al., 2024). Auto-tuning for Triton (Tillet et al., 2019) and Ladder (Wang et al., 2024) was enabled, while QuantLLM (Xia et al., 2024) used its heuristic policy to select kernel hyper-parameters. For end-to-end evaluations, we integrate our quantized kernels into the state-of-the-art LLM serving framework vLLM (Kwon et al., 2023) and compare them against vLLM (Kwon et al., 2023) and Ladder (Wang et al., 2024) in end to end execution. The specific versions of the tools are: vLLM v0.5.3, Triton v3.1.0, bitblas v0.0.1.dev15 (Ladder), QuantLLM with commit 9802c5a, and Marlin v0.1.1.
Hardware Configuration. Experiments were primarily conducted on a server equipped with an NVIDIA L40S GPU (48 GiB), with GPU driver 565.57.01 and CUDA Toolkit 12.6.3. Benchmarks were also performed on NVIDIA A100 and H100 GPUs to demonstrate the general applicability of our approach across different hardware platforms.
Experimental Protocol. For operator experiments, each kernel was executed $50$ times, while for model experiments, each model was executed $10$ times. In both cases, latency was measured using CUDA Events (NVIDIA Corporation, 2024a), and the median latency was reported. To eliminate artifacts from consecutive runs, the L2 cache was cleared before each execution.
8.2. Performance of Low-Precision Kernels
A single virtual machine program template is implemented to support matrix multiplication with all quantized types, taking tile sizes as tunable hyper-parameters. We denote the performance of this auto-tuned program as Tilus in the evaluation. Figure 9 compares the speedup of Triton (Tillet et al., 2019), Ladder (Wang et al., 2024), QuantLLM (Xia et al., 2024), Marlin (Frantar et al., 2024), and Tilus (ours) against cuBLAS (NVIDIA Corporation, 2023) for various low-precision matrix multiplications: uint8 (u8), float6_e3m2 (f6), uint4 (u4), int4 (i4), uint2 (u2), and uint1 (u1). While each baseline supports a limited set of quantized data types, Tilus consistently achieves speedups across all cases. For small batch sizes, the primary bottleneck is loading weights from global memory to registers for computation on SIMT or Tensor Cores. Triton struggles here due to costly layout conversions after weights are loaded into registers. Although preemptive conversion in global memory could mitigate this, Triton’s programming model lacks explicit layout control, making such optimizations infeasible. Ladder improves upon Triton by modifying data layouts in global memory, avoiding redundant conversions. However, it lacks critical optimizations such as software pipelining (NVIDIA Corporation, 2021; Huang et al., 2023), and its type-level packing limits efficient support for arbitrary bit widths, leading to underutilized memory bandwidth. Expert-crafted kernels from QuantLLM (Xia et al., 2024) and Marlin (Frantar et al., 2024) are optimized for specific quantization schemes but lack flexibility and maintainability. In contrast, Tilus outperforms all baselines with a single parameterized virtual machine template, efficiently supporting a full range of quantization types through a well-abstracted programming model.
8.3. Arbitrary Data Type Support
Figure 10. Speedup of quantized matrix multiplication compared against the cuBLAS FP16 kernel. A full spectrum of quantized data types is evaluated.
Our virtual machine (VM) supports low-precision matrix multiplications of the form matmul(A, B), where operand A can have data types with 32, 16, or 8 bits, and weight B supports a wide range of bit widths, from 32 bits down to 1 bit. Standard data types such as float32, float16, and int8 are supported, along with customized low-precision types with fewer than 8 bits, which include signed integers, unsigned integers, and floating-point formats with arbitrary exponent and mantissa distributions. Leveraging the algebraic layout system (Section 4 and 6.2), our VM enables efficient memory access for low-precision data. Figure 10 illustrates the speedup achieved for the full spectrum of quantized weight data type: uint1 to uint8, int2 to int8, and float3 to float8. Representative exponent-mantissa distribution of floating-point data types such as e4m3, e3m3, e3m2, e2m2, e2m1, and e1m1 are chosen. Each row represents the type kind (e.g., unsigned integer, signed integer or floating data type) while each columns represents the bit width. We take 128 as the group size of sub-channel quantization. Using matrix multiplication dimensions of BS=16, K=8192, and N=57344 the results demonstrate substantial speedups. These findings validate that our VM can effectively support arbitrary low-precision types with high efficiency, making it a robust solution for low-precision computations in modern GPUs. We want to highlight that all the kernels are from the same program template, by parameterizing tile sizes, thus the programming efforts are limited. There are around 200 configurations per operator, and it takes around one minute to compile. We used float16 as the activation data type in the experiment and we also support bfloat16 and int8.
8.4. End-to-End Performance
Figure 11. End-to-end performance across representative LLMs. The first two columns correspond to the decode stage with 1 and 16 tokens, respectively, while the third column shows latency for the prefill stage with 2048 prompt tokens.
We evaluated the end-to-end performance of representative LLMs: Gemma-2-9B (Team et al., 2024), QWen-2.5-32B (Yang et al., 2024), and Llama-3.3-70B (Grattafiori et al., 2024), across both prefill and decode stages. The prefill stage processes all prompt tokens at once, generating the kv-cache for subsequent token generation during the decode stage, which iteratively generates one token at a time. Prefill latency determines the time-to-first-token (TTFT), while decode latency impacts the speed of subsequent token generation. Both stages are critical for optimizing user experience and system utilization. Contiguous batching (Yu et al., 2022; Kwon et al., 2023) was used to batch multiple decode requests efficiently. Figure 11 shows the latency of both stages across these models. Our method consistently outperforms Ladder (Wang et al., 2024), particularly in the decode stage with more than one token (middle column of Figure 11). Analysis of Ladder’s generated kernels revealed suboptimal use of CUDA Cores for 1–15 tokens and Tensor Cores for 16 or more tokens, as key optimizations like software pipelining (Huang et al., 2023) and k-dimension parallelization (Osama et al., 2023) were not implemented, leading to poor performance. For the prefill stage, we decode quantized weights to float16 and perform computations using standard f16xf16 matrix multiplication kernels, as computation becomes the bottleneck at this stage. Our efficient handling of quantized weight layouts ensures minimal overhead for decoding, contributing to the superior performance observed.
8.5. Case Studies
8.5.1. Speedup over Different Hardware
Figure 12. End-to-end performance of the QWen2.5-30B model across NVIDIA A100, L40S, and H100 GPUs. The weight data types for vLLM, Ladder, and Tilus are float16, uint4, and uint4, respectively. OOM indicates out-of-memory error, and ERR indicates a runtime error.
We evaluate the end-to-end performance of the QWen2.5-30B model on NVIDIA A100, L40S, and H100 GPUs, which correspond to the Ampere, Ada Lovelace, and Hopper architectures, respectively. Figure 12 presents a performance comparison of vLLM (Kwon et al., 2023) (float16), Ladder (Wang et al., 2024) (uint4), and Tilus (uint4, ours) across the decode and prefill stages. On the Hopper architecture (H100), Ladder fails to generate valid kernels, leading to a CUDA error (’an illegal instruction was encountered’), which we denote as ERR in the figure. On the L40S GPU, vLLM (Kwon et al., 2023) exceeds the available 48 GiB DRAM capacity, leading to out-of-memory (OOM) errors. In all other configurations, Tilus consistently outperforms Ladder across all GPUs and both processing stages, highlighting its robust performance and adaptability across architectures.
8.5.2. Speedup over Different Batch Sizes
Figure 13. Speedup of quantized matmuls across different batch sizes from both prefill and decode stages.
We analyze the relationship between speedup and batch size by benchmarking matrix multiplication performance under different batch sizes. For the decode stage, we evaluate batch sizes of 1, 4, 8, and 16, while for the prefill stage, we use batch sizes of 4096, 8192, and 12,288. Experiments are conducted on Llama-3.3-70B-Instruct (Grattafiori et al., 2024) model with quantized data types float6_e3m2 (f6) and uint4 (u4), using $k=8192$ and $n=57344$. As shown in Figure 13, Tilus consistently outperforms baseline approaches across all batch sizes that are used in both decode and prefill stages of LLM serving. To further understand these performance differences, we profile kernels from cuBLAS, Ladder, and Tilus, with detailed analysis provided in Appendix D.
9. Related Work
Many deep learning compilers adopt loop-oriented scheduling (Chen et al., 2018; Ragan-Kelley et al., 2013) and build auto-tuning frameworks on top of it (Adams et al., 2019; Vaswani et al., 2017; Vasilache et al., 2018; Ahn et al., 2020; Xing et al., 2022; Zheng et al., 2022b, 2020a, a, 2020b; Shao et al., 2024; Feng et al., 2023; Wang et al., 2024). In contrast, Tilus employs a procedure-oriented approach that better models GPU hardware, improving programmability and flexibility. Beyond loop-oriented scheduling, tensor programs are often optimized using vendor libraries (e.g., cuBLAS (NVIDIA Corporation, 2023)), predefined templates for efficient matrix multiplication (NVIDIA Corporation, 2021), hardware-aware tiling strategies (Zhu et al., 2022), and domain-specific compilers for linear algebra (Sabne, 2020). While these methods prioritize performance, they lack extensibility for arbitrary low-precision data types. Other research focuses on optimizing irregular or ragged tensor programs (Tang et al., 2022; Fegade et al., 2022), operator fusion (Zheng et al., 2022c; Zhao et al., 2022), dynamic shape handling (Shen et al., 2021; Zhu et al., 2021; Fegade et al., 2021; Zheng et al., 2022b), and scheduling independent operators (Ma et al., 2020; Ding et al., 2021; Jia et al., 2019). Microscaling data types (Rouhani et al., 2023) can be thought as a more fine-grained quantization thus we could also support it. These techniques are complementary to our focus on efficient low-precision computation. Triton (Tillet et al., 2019) introduces a tile-based programming model. However, it lacks explicit support for low-precision data types and does not expose the GPU memory hierarchy, limiting optimization opportunities. Similarly, Hidet (Ding et al., 2023), which serves as our backend, does not provide built-in support for low-precision types. Graphene (Hagedorn et al., 2023) presents an intermediate representation (IR) with a layout representation. Unlike Graphene’s focus on strides and computation, our algebraic layout system emphasizes hierarchical organization. In fact, we can express Graphene’s layout representation as a primitive component within our system.
10. Conclusion
We introduced Tilus, a GPGPU virtual machine designed for efficient low-precision LLM serving, addressing key limitations of existing solutions. Tilus features an algebraic layout system for tensor distribution within thread block registers, a thread-block-level programming model with fine-grained memory management, and broad support for sub-byte data types, enabling arbitrary precision from 1 to 8 bits. Experimental results show substantial performance gains over state-of-the-art frameworks like Triton and Ladder, demonstrating the flexibility and scalability of our approach. This work establishes a foundation for efficient and extensible LLM inference, paving the way for further optimizations in emerging hardware, advanced quantization techniques, and diverse low-precision formats.
Appendix
Appendix A Layout Formalization
In this appendix section, we formalize the concept of layout, define primitive layouts, composition, division operators, and layout broadcast. Furthermore, we prove that the set of layouts with the composition operator forms a monoid.
Definition 0 (Layout).
A layout $f$ is a function that maps a pair of thread index $t$ and local element index $i$ to the logical index $I$ of the element in the tensor:
[…]
where $T$ is the number of threads, $m$ is the number of elements each thread holds, $r$ is the rank of the tensor, $\mathbb{N}_{n}=\{0,1,\dots,n-1\}$ and $S=(n_{1},n_{2},\dots,n_{r})$ is the shape of the tensor.
Definition 0 (Local Layout).
Given an $r$-rank tensor with shape $(n_{1},n_{2},\dots,n_{r})$, we define the local layout as a layout with a single thread ($T=1$), where all elements of the tensor are stored in row-major order:
$f(t,i)=(i/(n_{2}n_{3}\dots n_{r}),\;i/(n_{3}n_{4}\dots n_{r})\%n_{2},\;\dots,\;i\%n_{r}).$
Here, $0\leq t<1=T$ and $0\leq i<n_{1}n_{2}\dots n_{r}=m$. In code, we represent such a layout as local(n1, n2, ..., nr).
Definition 0 (Spatial Layout).
Given an $r$-rank tensor with shape $(n_{1},n_{2},\dots,n_{r})$, we define the spatial layout as a layout that distributes all elements across $n_{1}n_{2}\dots n_{r}$ threads, where each thread holds a single element:
$f(t,i)=(t/(n_{2}n_{3}\dots n_{r}),\;t/(n_{3}n_{4}\dots n_{r})\%n_{2},\;\dots,\;t\%n_{r}).$
Here, $0\leq t<n_{1}n_{2}\dots n_{r}=T$ and $0\leq i<1=m$. In code, we represent such a layout as spatial(n1, n2, ..., nr).
Definition 0 (Layout Composition).
Given two layouts $f$ and $g$ with the same rank $r$, we define the composition layout $h$ of $f$ and $g$ as:
$h(t,i)$ $=f(t\;/\;T_{g},i\;/\;m_{g})\odot S_{g}+g(t\;\%\;T_{g},i\;\%\;m_{g}),$
$T_{h}$ $=T_{f}T_{g},$
$m_{h}$ $=m_{f}m_{g},$
$S_{h}$ $=S_{f}\odot S_{g},$
where $T_{h,f,g}$, $m_{h,f,g}$, and $S_{h,f,g}$ are the number of threads, local elements per thread, and shape of layouts $h,f,g$, respectively. Here, $\odot$ represents elementwise multiplication. The composition is denoted as $h=f\circ g$. In code, we use spatial(2, 3).local(3, 2) to denote the composition of spatial(2, 3) and local(3, 2).
Theorem 5 (Layout Composition Associativity).
The layout composition operation is associative:
$f\circ(g\circ h)=(f\circ g)\circ h.$
Proof.
Let $f,g,h$ be layouts with the following properties:
$f(t,i):\mathbb{N}_{T_{f}}\times\mathbb{N}_{m_{f}}\to\mathbb{N}_{n_{1}}\times\cdots\times\mathbb{N}_{n_{r}},$
$g(t,i):\mathbb{N}_{T_{g}}\times\mathbb{N}_{m_{g}}\to\mathbb{N}_{n_{1}}\times\cdots\times\mathbb{N}_{n_{r}},$
$h(t,i):\mathbb{N}_{T_{h}}\times\mathbb{N}_{m_{h}}\to\mathbb{N}_{n_{1}}\times\cdots\times\mathbb{N}_{n_{r}}.$
For $f\circ g$, the layout $k(t,i)$ is:
$k(t,i)=f(t/T_{g},i/m_{g})\odot S_{g}+g(t\%T_{g},i\%m_{g}),$
where $T_{k}=T_{f}T_{g},\;m_{k}=m_{f}m_{g},\;S_{k}=S_{f}\circ S_{g}$.
For $g\circ h$, the layout $l(t,i)$ is:
$l(t,i)=g(t/T_{h},i/m_{h})\odot S_{h}+h(t\%T_{h},i\%m_{h}),$
where $T_{l}=T_{g}T_{h},\;m_{l}=m_{g}m_{h},\;S_{l}=S_{g}\circ S_{h}$.
Now consider $f\circ(g\circ h)$. we get layout $p$:
$p(t,i)=f(t/T_{l},i/m_{l})\odot S_{l}+l(t\%T_{l},i\%m_{l}).$
Expanding $l(t\%T_{l},i\%m_{l})$:
$l(t\%T_{l},i\%m_{l})=$ $g((t\%T_{l})/T_{h},(i\%m_{l})/m_{h})\odot S_{h}$
$+h((t\%T_{l})\%T_{h},(i\%m_{l})\%m_{h}).$
Substituting this into $p(t,i)$:
$p(t,i)=$ $f(t/T_{l},i/m_{l})\odot S_{l}$
$+g((t\%T_{l})/T_{h},(i\%m_{l})/m_{h})\odot S_{h}$
$+h((t\%T_{l})\%T_{h},(i\%m_{l})\%m_{h}).$
For $(f\circ g)\circ h$, we get layout $q$:
$q(t,i)=k(t/T_{h},i/m_{h})\odot S_{h}+h(t\%T_{h},i\%m_{h}).$
Expanding $k(t/T_{h},i/m_{h})$:
$k(t/T_{h},i/m_{h})=$ $f((t/T_{h})/T_{g},(i/m_{h})/m_{g})\odot S_{g}$
$+g((t/T_{h})\%T_{g},(i/m_{h})\%m_{g}).$
Substituting this into $q(t,i)$:
$q(t,i)=$ $f(t/T_{l},i/m_{l})\odot S_{l}$
$+g((t\%T_{l})/T_{h},(i\%m_{l})/m_{h})\odot S_{h}$
$+h((t\%T_{l})\%T_{h},(i\%m_{l})\%m_{h}).$
Since $p(t,i)=q(t,i)$, we conclude:
$f\circ(g\circ h)=(f\circ g)\circ h.$
This completes the proof. ∎
Definition 0 (Layout Broadcast).
To handle layouts of different ranks, we define the layout broadcast operation, which increases the rank of a layout by prepending zeros to the output.
Let $f(t,i):\mathbb{N}_{T}\times\mathbb{N}_{m}\to\mathbb{N}_{n_{1}}\times\cdots\times\mathbb{N}_{n_{r}}$ be a layout of rank $r$. To broadcast $f$ to rank $r^{\prime}>r$, we define:
$f^{\prime}(t,i)=(0,\dots,0,f(t,i)),$
where $r^{\prime}-r$ zeros are prepended. In code, this is denoted as broadcast(f, rank).
Using broadcast, we can compose any two layouts $f$ and $g$ of ranks $r_{f}$ and $r_{g}$, respectively, by broadcasting the smaller rank to match the larger one before composition:
$f\circ g=\text{broadcast}(f,\max(r_{f},r_{g}))\circ\text{broadcast}(g,\max(r_{f},r_{g})).$
Definition 0 (Layout Division).
Given two layouts $f$ and $g$, if there exists a layout $h$ such that $f=h\circ g$, we define $h$ as the result of dividing $f$ by $g$:
$h=f/g.$
In code, this is denoted as f/g.
Figure 14. The optimized matrix multiplication with proper tile sizes and software pipeline optimization.
Theorem 8 (Monoid Structure of Layouts).
The set of layouts $L$, under the composition operator $\circ$, forms a monoid.
Proof.
To prove that $(L,\circ)$ is a monoid, we verify the three properties:
  • Clousure: The composition of two layouts is a layout, as shown in the definition of composition.
  • Associativity: This was proven earlier in the theorem on layout composition associativity.
  • Identity Element: Define the identity layout $e$ as local(1) (or equivalently spatial(1)). It satisfies $f\circ e=e\circ f=f$ for any layout $f$.
Thus, $(L,\circ)$ satisfies the monoid properties. ∎
Appendix B Optimized Matrix Multiplication
In Figure 2 of the main paper, we presented an example of quantized matrix multiplication (FP16xINT6). This example primarily demonstrates the programming of the virtual machine using algebraic layouts, thread-block-level instructions, and the handling of low-precision data types. However, it does not incorporate optimized tiling sizes or advanced techniques such as software pipelining (Huang et al., 2023).
In this section, we extend the example from the main paper by introducing software pipelining and selecting better tiling sizes to program an efficient matrix multiplication. The complete program is illustrated in Figure 14.
B.1. New Optimizations
The key improvement in this program compared to the simplified version in Figure 2 is the use of shared memory to store global sub-tensors, which significantly enhances data reuse since shared memory is an order of magnitude faster than global memory (NVIDIA Corporation, 2024a). Shared memory is also employed as a staging area for results before writing them back to global memory, ensuring coalesced global memory access (NVIDIA Corporation, 2024a). Asynchronous data transfer from global memory to shared memory is achieved using the CopyAsync, CopyAsyncCommitGroup, and CopyAsyncWaitGroup instructions. To maximize performance, memory loading is pipelined with computation. This is accomplished by allocating three shared memory buffers for each operand (A and B). During the main loop (Line 31-45), the shared memory buffer for the next stage (preload_stage) is populated (Line 39-40) while computations are performed using the buffer for the current stage (current_stage). These optimizations are essential to achieve state-of-the-art performance. Our virtual machine not only supports these optimizations seamlessly but also accommodates arbitrary low-bit-width quantized data types via register tensor reinterpretation (Line 36).
B.2. Line-by-Line Code Analysis
Initialization: Workload sizes, tiling sizes, and the number of pipeline stages are defined (Line 1-3). Register and shared memory tensor layouts for the virtual machine program are then specified (Line 5-11). There are two new operations over the layouts that we did not introduce in the main text: reduce(...) and swizzle(...). The reduce(...) layout reduces a layout over some dimensions. If there are multiple indices over that dimension are stored in the same thread, the thread will only keep one copy of the elements. If there are multiple threads along the reduction dimension, we will retain one copy of the element for each thread. In this example, reduce(spatial(1, 1, 4), dims=[2]) is a layout with shape $(1,1)$ and $4$ threads. Each thread will hold the element $(0,0)$. On the other hand, the swizzle operator swizzle a layout by performing the following operation:
$\text{swizzle}(i,j)=$ $\left(i,j\oplus\left\lfloor\frac{i}{2^{\text{log\_step}}}\right\rfloor\right)$
$g(t,i)=$ $\text{swizzle}(f(t,i))$
where $\oplus$ is bitwise-xor operation, and log_step is the parameter of the swizzle operator. The dim=1 in the code indicates that we hope to swizzle the dimension $1$, that’s why we changed the dimension $1$ and keep the other dimension unchanged. The swizzle operation is necessary to avoid the shared memory bank-conflict (NVIDIA Corporation, 2024a) for both the global to shared memory transfer (via CopyAsync) and the shared memory to register transfer (via LoadShared).
Global Tensor Views and Memory Allocation: Global tensor views and offsets for thread blocks are initialized (Line 14-18). Shared memory for software pipelining and registers for accumulation are allocated (Line 19-21).
Preloading Shared Memory: The first STAGES - 1 iterations preload the shared memory buffers (Line 22-28). Two variables, current_stage and preload_stage, track the buffers used for computation and memory copying (Line 29-30).
Main Loop: For each iteration, the k-dimension is further tiled to reduce register usage (Line 33-38). Shared memory buffers are loaded into register tensors, reinterpreted for quantized data types, and cast to the computation data type (e.g., FP16). The dot product operation is then performed. Simultaneously, asynchronous global-to-shared memory transfers are issued using a separate buffer (Line 39-44). This multi-buffer approach enables parallel utilization of GPU computation and memory transfer units.
Result Storing: After completing the main loop, the accumulation tensor is cast from float32 to float16 (Line 46). The tensor is stored in shared memory (Line 48), reloaded into a new register tensor (Line 50), and written back to global memory with coalesced access (Line 51).
This improved program illustrates the flexibility of the virtual machine in implementing advanced optimizations while supporting low-precision data types efficiently. By utilizing shared memory, asynchronous transfers, and software pipelining, it achieves significant performance gains.
Appendix C Program Compilation and Runtime
We provides a domain-specific language (DSL) for our virtual machine in Python to allow programmers to directly write the virtual machine programs in Python, making it easy to integrate the produced kernels with the rich deep learning ecosystem in Python. Given a program, we take several steps to compile it to GPU executable code.
Step 1: Global and Shared Memory Planning Each GPU kernel could use some launch-time known size of shared memory space. To simplify GPU programming, we allow the users to allocate shared memory multiple times in the program on demand. Thus, we need a shared memory planner to calculate the size of shared memory the virtual machine needs, and map the shared tensor to one region of the kernel’s shared memory space. Similar to shared memory planning, we also require a global memory planner to manage the allocation of global memory shared by all thread blocks. We will request the runtime system of our virtual machine to allocate a workspace in global memory, enabling the kernel to use this workspace during its execution.
Step 3: Code Emitting for Each Instruction. We emit low-level GPU code for each VM instruction one by one. In our implementation, we use the Hidet IR (Ding et al., 2023) to represent the low-level GPU code. During this process, we will do instruction selection to select the most efficient low-level instructions when feasible. For example, we use lds PTX instruction (NVIDIA Corporation, 2024b) to load the data from shared memory to register. However, a more efficient PTX instruction ldmatrix could also be used if the layout of the loaded register tensor is divisible by layout spatial(8, 4).repeat(1, 4). Besides, we also try to perform automatic vectorization for memory loading and storing instructions. For example, we will try to use the vectorized instructions like cp.async.v4, lds128 and ldg128 to maximize the memory accessing efficiency.
Step 4: Lowering Low-Precision Data Types. After we emit thread-block-level instructions to the low-level IR, we will apply the passes that implement the the rules discussed in Section 7 to transform all low-precision operations in the low-level IR to corresponding operations on hardware-friendly types. en most cases, only the vectorized type casting from low-precision type to standard type (e.g., float16) will be applied since the memory loading of low-precision data will be replaced by standard types thanks to our layout formalization and register tensor reinterpretation. Afterwards, we generate the CUDA (for NVIDIA GPUs) code from the low-level IR, and finally use the compiler nvcc to compile the source code to hardware binary.
Vitual Machine Runtime. The compiled binary could be loaded by the virtual machine runtime. The runtime also maintains internal states to serve the kernel execution: 1) an on-demand allocated workspace memory that can be requested by the compiled kernels; 2) an execution context that stores the CUDA stream that the kernel will be launched on; 3) the kernels cached in memory.
Appendix D Profiling of Quantized Kernels
Figure 15. GPU hardware utilization for cuBLAS, Ladder, and Tilus kernels. The figure details the throughput utilization of memory units and pipeline utilization of computation units. Nsight Compute is used to collect the results.
We use Nsight Compute to profile the kernels of cuBLAS, Ladder, and Tilus, with results shown in Figure 15. The figure details GPU hardware utilization, dividing it into throughput utilization for L1 Cache, L2 Cache, and DRAM (first three columns) and computation unit utilization for ALU, FMA, and Tensor Core (last three columns). Specifically, ALU handles bit manipulation and logical instructions, FMA performs fused multiply-add operations for integers and floating-point numbers, and Tensor Core represents the utilization of NVIDIA’s specialized Tensor Core pipeline (NVIDIA Corporation, 2024a). From row B of Figure 15, Ladder kernels primarily rely on CUDA Cores, with $40.5\%$ FMA utilization and no Tensor Core usage. However, the kernel suffers from significant DRAM bandwidth limitations ($96.6\%$ utilization), as each thread independently loads operand A from global memory without exploiting shared memory for data reuse. This inefficiency worsens with larger batch sizes, as seen in Figure 13, where performance degrades for batch sizes of 1, 4, and 8 (Wang et al., 2024). For batch sizes of 16 or more, Ladder transitions to Tensor Core usage (row C, Figure 15). Despite this transition, Ladder does not utilize cp.async for asynchronous memory transfers or software pipelining to overlap computation and memory operations. These limitations hinder efficient resource utilization. Additionally, Ladder’s Tensor Core execution strategy is flawed; padding in the batch size dimension results in tensor core instructions up to eight times more than necessary. In contrast, Tilus overcomes these inefficiencies using an algebraic layout system. It interprets quantized weights with standard data types and employs cp.async to optimize memory access patterns. Moreover, Tilus integrates software pipelining to overlap memory transfers and computation, leading to superior performance, as evident in row D of Figure 15.